OpenStudy (anonymous):

\[\int\limits_{0}^{3}\int\limits_{0}^{\sqrt{(9-x^2)}}\int\limits_{0}^{3-x} (xz + 3z) dy dz dx\]

OpenStudy (anonymous):

how to integrate?

OpenStudy (anonymous):

Not 100% sure but to me it just looks like simple integration. All you have to do is integrate in respect to y first then use the limits 0 to 3-x. Then whatever you get you integrate in respect to z and then you use the limits of 0 to sqrt( 9-x^2) then integrate whatever you get in respect to x using the limits of 0 to 3 The only big problem you will have is integrating in respect to x since that will be messy

OpenStudy (anonymous):

\[\int\limits_{0}^{3}\int\limits_{0}^{\sqrt{(9-x^2)}}\int\limits_{0}^{3-x} (xz + 3z) dy dz dx\]